Supply Demand 10 20 30 40 50 60 70 80 Quantity

  1. How much will the buyer pay for the product after the tax is imposed?
  2. How much will the seller receive after the tax is imposed?
  3. As a result of the tax, what has happened to the level of output?
  4. Calculate the economic welfare after government imposes a tax of $5 per unit on buyers.
    1. Total Surplus
    1. Government Revenue
    1. DWL
    1. Producer Surplus
Answer #1

(1)The price paid by buyer for the product after the imposition of tax is =$10 (clearly Shown in the diagram)

(2) The seller receives the price for the product after the imposition of tax is =$5 (clearly Shown in the diagram)

(3) Before the tax was imposed the equilibrium level of output is 40 units because at 40 units of quantity of output the demand curve equals with the supply curve and the equilibrium price was $8.

                Due to imposition of tax the new equilibrium level of output is 20 units, which is less than the before tax equilibrium quantity

(4) The economic welfare after government imposes tax on consumer or consumer surplus

Because consumer surplus deals with the surplus obtained by buyer

Demand choke price = $12

Price paid by the buyer for the product is =$10

Equilibrium level of output after tax is 20 units

Consumer surplus = 0.5x(12-10)x(20-0) = 20

The economic welfare obtained by buyer after government imposes tax = 20

(5) Producer surplus is concerned with supplier

Supply choke price = $2

The price received by the seller for the product after the imposition of tax is =$5

Equilibrium level of output after tax is 20 units

Producer surplus = 0.5x(5-2)x(20-0) = 30

Producer surplus = 30

(6) Total welfare or Total surplus = Consumer surplus + producer surplus = 20 +30 =50

Total welfare or Total surplus = 50

(7) Dead weight loss (DWL)

Price paid by the buyer for the product after tax is =$10

The price received by the seller for the product after the imposition of tax is =$5

Equilibrium level of output after tax is 20 units

Equilibrium level of output before tax is 40 units

Dead weight loss = 0.5x(10-5)x(40-20)=50

Dead weight loss (DWL) is = 50

(8) Government revenue

Price paid by the buyer for the product after tax is =$10

The price received by the seller for the product after the imposition of tax is =$5

Equilibrium level of output after tax is 20 units

Government revenue = (10-5)x(20-0) = 100

Government revenue =100